# Shot-noise in strongly-correlated double quantum spin Hall edges

### v2

@@ Abstract
To model a QPC, we only consider scattering between the top and bottom edges at a single point. Away from this point, the edges are taken to be fully decoupled. As we shall now discuss, we only need to consider processes where one or two electrons are scattered across the QPC. Let us consider $I_{\text{b},m}$, the contribution to the average backscattered current from a process where $m$-particles are backscattered. We will consider the low voltage limit, where the applied voltage $V$ is small compared to the characteristic energy scale of the wire, $v_{\text{F}}/L,$ which is the ratio of the plasmon frequency over the length of the wire. In this limit, $I_{\text{b},m}$ scales as $V^{m^{2}(g_{\infty}+1/g_{\infty})-1}$, where $g_{\infty}$ is the value of the Luttinger parameter near the leads [55, 33, 54]. For Fermi-liquid-like leads, $g_{\infty}=1$, independent of the value of the Luttinger parameter in the bulk. At sufficiently low voltage, the shot noise signal is therefore dominated by processes with the lowest number of backscattering particles. For the weakly-correlated edge, the lowest number of particles that can be backscattered is $m=1$. For the strongly-correlated edge, there is a gap to one-electron scattering, and so the dominant contribution to the noise occurs from $m=2$-electron scattering.
IV.2.1 Backscattering of the weakly-correlated edge.
First, we will consider the backscattering action for the weakly-correlated edge. The symmetry-allowed backscattering operators are
@@
In bosonized form, the backscattering action for the weakly-correlated edge is (see Appendix B)
@@
IV.2.2 Backscattering of the strongly-correlated edge.
For the strongly-correlated edge, there is a gap to single-electron operators. However, there are still relevant two-electron tunneling operators at the edge,
@@
The first term describes tunneling of the charge $2e$ spin 00 modes in Eq. 22, while the second term describes the tunneling of the charge $2e$ spin ±$\pm 1$ modes in Eq. 21. Again, all fields are evaluated at $\bm{x}=(t,x_{0})$, and $\lambda^{\prime}_{i}$ are the strengths of the different tunneling processes. The difference in the coefficient of $\varphi_{\text{c}}$ in the cosine terms in Eqs. 34 and 35 is ultimately responsible for the difference in the Fano factors, as we will show in the following section.
@@
We now calculate the Fano factor for the weakly-correlated edge, and the strongly-correlated edge. Since the charge is carried by the $\varphi_{\text{c}}$ boson for both types of edges, the analysis of the weakly-correlated edge and the strongly-correlated edge is very similar in our current formulation.
IV.3.1 Weakly-correlated edge.
Our starting point is the action $S=S_{0}+S_{\text{b}}+S_{V}$, where $S_{0}$ and $S_{\text{b}}$ are the bare and backscattering action of the weakly-correlated edge discussed previously, and $S_{V}$ describes the coupling to the external leads of the system, which we will define below. In the folded picture, we assume the leads are at $x=\pm L/2$, where $L$ is the length of the edge, and that the QPC is somwhere between the leads, $-L/2<x_{0}<L/2$.
When there is a voltage applied across the system, the edge is brought out of equilibrium. Thus, we must consider time evolution along the Keldysh contour [56], which we denote by $\Gamma$. This contour takes an excursion from $\infty$- ∞ to $\infty$ above the real time axis, and then returns to $\infty$- ∞ below it. In this formulation, the charge and spin part of the quadratic action is
$\mathcal{L}_{0,\text{cs}}=\bm{\mathbf{\varphi}}_{\text{c}}^{\intercal}\hat{\bm{\mathbf{D}}}_{\text{c}}\bm{\mathbf{\varphi}}_{\text{c}}+\bm{\mathbf{\varphi}}_{\text{s}}^{\intercal}\hat{\bm{\mathbf{D}}}_{\text{s}}\bm{\mathbf{\varphi}}_{\text{s}}$ (36)
[…]
$\hat{\bm{\mathbf{D}}}_{\gamma}^{\alpha,\beta}=\left[G_{\gamma}^{\alpha,\beta}\right]^{-1}$ (37)
@@
IV.3.2 Shot noise of the strongly-correlated edge.
We now consider shot noise for the strongly-correlated edge. The analysis is largely the same as for the weakly-correlated edge, as the current is still carried by the $\varphi_{\text{c}}$ boson in the strongly-correlated edge. The steps from Eqs. 39-45 are unchanged. However, instead of Eq. 46 we have (see Appendix F)
@@
In this work, we analyzed the role of interactions at the edge of the double quantum spin Hall insulators realized in twisted TMDs. We have shown that the edge has two symmetry preserving phases: a weakly-correlated phase, a strongly-correlated phase. The weakly-correlated edge has two pairs of helical modes and is adiabatically connected to two decoupled quantum spin Hall edges. The strongly-correlated edge has a single helical mode, and notably, is gapped to single electrons. In contrast, pairs of electrons remain gapless, leading the strongly-correlated edge to have a Fano factor of $\mathcal{F}=2$ in shot-noise measurements. Since $\mathcal{F}=1$ for the weakly-correlated edge, shot noise measurements are a concrete method for distinguishing between the two possible edges.
While we have only considered the double quantum spin Hall insulator, it is straightforward to generalize these results to $N$-fold QSHIs. In this case, there is a non-trivial strongly-correlated edge where there is a gap to single electrons, and the local degrees of freedom carry charge $Ne$. The Fano factor of the strongly-correlated $N$-fold QSHI edge will therefore be $\mathcal{F}=N$. In particular, if the experimentally-observed triple quantum spin Hall insulator [16, 18] realizes a strongly-correlated edge, the Fano factor will be $\mathcal{F}=3$.
In general, the existence of the strongly-correlated edge depends on the Luttinger parameter of the edge theory. The Luttinger parameter, in turn, depends on the velocity of the edge modes with a lower bare velocity producing a smaller Luttinger parameter and stronger correlations. There are two main factors that determine the bare velocity of the edge modes: the properties of the bulk band dispersion and the microscopic details of the edge termination. The bulk bands of twisted TMDs are quite flat, making them potential candidates for realizing the strongly-correlated DQSHI edge. However, the physical boundaries of the twisted TMD must still be taken into account to confirm this expectation. To measure the doubled Fano factor of the strongly-correlated edge, it is also important that the constriction used to create the QPC does not destabilize the strongly-correlated edge itself. If the single electron gap vanishes at the QPC, then individual electrons can tunnel between the edges, as in the weakly-correlated edge. More sophisticated modeling of the boundaries and constrictions of a realistic TMD-device is needed to determine if if the effects discussed here will occur or not.
@@ Acknowledgments.
$\bm{\mathbf{\varphi}}^{\prime}\equiv\bm{\mathbf{W}}^{-1}\bm{\mathbf{\varphi}}$ (50)
@@
$\bm{\mathbf{W}}=\frac{1}{2}\begin{pmatrix}1&1&1&1\\ -1&-1&1&1\\ 1&-1&-1&1\\ -1&1&-1&1\end{pmatrix},\text{ and }\bm{\mathbf{W}}^{-1}=\bm{\mathbf{W}}^{\intercal}=\frac{1}{2}\begin{pmatrix}1&-1&1&-1\\ 1&-1&-1&1\\ 1&1&-1&-1\\ 1&1&1&1\end{pmatrix}$ (51)
@@
$\bm{\mathbf{\varphi^{\prime}}}=\begin{pmatrix}\varphi_{+}\\ \varphi_{-}\\ \theta_{-}\\ \theta_{+}\end{pmatrix}.$ (52)
@@
$\begin{split}\mathcal{L}&=-\frac{1}{4\pi}\partial_{x}(\bm{\mathbf{W}}\bm{\mathbf{\varphi}}^{\prime})^{\intercal}\bm{\mathbf{K}}\partial_{t}(\bm{\mathbf{W}}\bm{\mathbf{\varphi}}^{\prime})+\frac{1}{4\pi}\partial_{x}(\bm{\mathbf{W}}\bm{\mathbf{\varphi}}^{\prime})^{\intercal}\bm{\mathbf{V}}\partial_{x}(\bm{\mathbf{W}}\bm{\mathbf{\varphi}}^{\prime})\\ &=-\frac{1}{4\pi}\partial_{x}\bm{\mathbf{\varphi}}^{\prime\intercal}\underbrace{(\bm{\mathbf{W}}^{\intercal}\bm{\mathbf{K}}\bm{\mathbf{W}})}_{\mathclap{\bm{\mathbf{K}}^{\prime}}}\partial_{t}\bm{\mathbf{\varphi}}^{\prime}+\frac{1}{4\pi}\partial_{x}\bm{\mathbf{\varphi}}^{\prime\intercal}\underbrace{(\bm{\mathbf{W}}^{\intercal}\bm{\mathbf{V}}\bm{\mathbf{W}})}_{\mathclap{\bm{\mathbf{V}}^{\prime}}}\partial_{x}\bm{\mathbf{\varphi}}^{\prime}\end{split}$ (53)
where the transformed matrix $\bm{\mathbf{K}}^{\prime}$ is now
$\bm{\mathbf{K}}=\begin{pmatrix}0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\\ 1&0&0&0\\ \end{pmatrix}$ (54)
@@
$\bm{\mathbf{V}}^{\prime}=\bm{\mathbf{W}}^{\intercal}\bm{\mathbf{V}}\bm{\mathbf{W}}=\begin{pmatrix}v-v^{\prime}+u-u^{\prime}&&&\\ &v-v^{\prime}-u+u^{\prime}&&\\ &&v+v^{\prime}-u-u^{\prime}&\\ &&&v+v^{\prime}+u+u^{\prime}\end{pmatrix}.$ (55)
@@
$\begin{split}\mathcal{L}=&\frac{1}{2\pi}\left[(\partial_{x}\varphi_{+})(\partial_{t}\vartheta_{+})+(\partial_{x}\varphi_{-})(\partial_{t}\vartheta_{-})\right]\\ &+\frac{1}{4\pi}\left[(v+v^{\prime}-u-u^{\prime})(\partial_{x}\vartheta_{-})^{2}+(v-v^{\prime}-u+u^{\prime})(\partial_{x}\varphi_{-})^{2}\right]\\ &+\frac{1}{4\pi}\left[(v+v^{\prime}+u+u^{\prime})(\partial_{x}\vartheta_{+})^{2}+(v-v^{\prime}+u-u^{\prime})(\partial_{x}\varphi_{+})^{2}\right]\\ \end{split}$ (56)
@@
$\bm{\mathbf{V}}=\begin{pmatrix}v_{1}&-v_{1}^{\prime}&u&-u^{\prime}\\ -v_{1}^{\prime}&v_{1}&-u^{\prime}&u\\ u&-u^{\prime}&v_{2}&-v_{2}^{\prime}\\ -u^{\prime}&u&-v_{2}^{\prime}&v_{2}\end{pmatrix}$ (58)
@@
where $w=v_{1}-v_{2}-v_{1}^{\prime}+v_{2}^{\prime}$ and $w^{\prime}=v_{1}-v_{2}+v_{1}^{\prime}-v_{2}^{\prime}$. However, we show in Appendix C that including these interactions do not affect the main results, since the propagators between the $+$ and -- sectors vanish in the limit as $\omega\to 0$.
@@
[…]
@@ Appendix C Low-energy asymptotics of the retarded propagator
In this section, we slightly generalize the calculation found in [51] to the case of two coupled inhomogeneous Luttinger liquids with unequal velocities and scattering parameters. We will obtain the low-energy asymptotic form of the various correlators, which we show are determined by the values of the parameters in the leads. We demonstrate that the original result of [51] still holds, and that the propagators between the $+$ and -- sectors vanish as $\omega\to 0$. This means that we may ignore this interaction term for the purposes of this analysis, which is concerned exclusively with the zero-frequency limit of all quantities.
@@
$\begin{split}\hat{\bm{\mathbf{D}}}\bm{\mathbf{G}}&=\begin{pmatrix}D_{+}&D^{\prime}\\ D^{\prime}&D_{-}\end{pmatrix}\begin{pmatrix}G_{+}&G^{\prime}\\ G^{\prime\prime}&G_{-}\end{pmatrix}\\ &=\delta(x-x^{\prime})\mathds{1}_{2\times 2}\\ \end{split}$ (69)
The entries of the matrix $\hat{\bm{\mathbf{D}}}$ are the following differential operators:
@@ C.3 Region II: $0<x<x^{\prime}$
By acting with the appropriate differential operators on the equations (for instance, $D^{\prime}$ on (85) and with $D_{-}$ on (87)), one finds that $G_{\pm}$, $G^{\prime}$ and $G^{\prime\prime}$ are all solutions to the same homogeneous, fourth-order, linear differential equation:
@@
$D_{-}D_{+}=\frac{v_{-}v_{+}}{K_{-}K_{+}}\partial_{x}^{4}-\omega^{2}\left(\frac{v_{+}^{2}+v_{-}^{2}}{v_{-}K_{-}v_{+}K_{+}}\right)+\frac{\omega^{4}}{v_{-}K_{-}v_{+}K_{+}}\$ (90)

### v3

@@ Abstract
To model a QPC, we only consider scattering between the top and bottom edges at a single point. Away from this point, the edges are taken to be fully decoupled. As we shall now discuss, we only need to consider processes where one or two electrons are scattered across the QPC. Let us consider $I_{\text{b},m}$, the contribution to the average backscattered current from a process where $m$-particles are backscattered. We will consider the low voltage limit, where the applied voltage $V$ is small compared to the characteristic energy scale of the wire, $v_{\text{F}}/L,$ which is the ratio of the plasmon frequency over the length of the wire. In this limit, $I_{\text{b},m}$ scales as $V^{m^{2}(g_{\infty}+1/g_{\infty})-1}$, where $g_{\infty}$ is the value of the Luttinger parameter near the leads [72, 43, 71]. For Fermi-liquid-like leads, $g_{\infty}=1$, independent of the value of the Luttinger parameter in the bulk. At sufficiently low voltage, the shot noise signal is therefore dominated by processes with the lowest number of backscattering particles. For the weakly correlated edge, the lowest number of particles that can be backscattered is $m=1$. For the strongly correlated edge, there is a gap to one-electron scattering, and so the dominant contribution to the noise occurs from $m=2$-electron scattering.
IV.2.1 Backscattering of the weakly correlated edge.
First, we will consider the backscattering action for the weakly correlated edge. The symmetry-allowed backscattering operators are
@@
In bosonized form, the backscattering action for the weakly correlated edge is (see Appendix B)
@@
IV.2.2 Backscattering of the strongly correlated edge.
For the strongly correlated edge, there is a gap to single-electron operators. However, there are still relevant two-electron tunneling operators at the edge,
@@
The first term describes tunneling of the charge $2e$ spin 0 modes in Eq. 25, while the second term describes the tunneling of the charge $2e$ spin $\pm 1$ modes in Eq. 24. Again, all fields are evaluated at $\bm{x}=(t,x_{0})$, and $\lambda^{\prime}_{i}$ are the strengths of the different tunneling processes. The difference in the coefficient of $\varphi_{\text{c}}$ in the cosine terms in Eqs. 37 and 38 is ultimately responsible for the difference in the Fano factors, as we will show in the following section.
@@
We now calculate the Fano factor for the weakly correlated edge, and the strongly correlated edge. Since the charge is carried by the $\varphi_{\text{c}}$ boson for both types of edges, the analysis of the weakly correlated edge and the strongly correlated edge is very similar in our current formulation.
IV.3.1 Shot noise of the weakly correlated edge.
Our starting point is the action $S=S_{0}+S_{\text{b}}+S_{V}$, where $S_{0}$ and $S_{\text{b}}$ are the bare and backscattering action of the weakly correlated edge discussed previously, and $S_{V}$ describes the coupling to the external leads of the system, which we will define below. In the folded picture, we assume the leads are at $x=\pm L/2$, where $L$ is the length of the edge, and that the QPC is somwhere between the leads, $L/2<x_{0}<L/2$.
When there is a voltage applied across the system, the edge is brought out of equilibrium. Thus, we must consider time evolution along the Keldysh contour [73], which we denote by $\Gamma$. This contour takes an excursion from -$\infty$ to $\infty$ above the real time axis, and then returns to -$\infty$ below it. In this formulation, the charge and spin part of the quadratic action is
$\mathcal{L}_{0,\text{cs}}=\boldsymbol{\mathbf{\varphi}}_{\text{c}}^{\intercal}\hat{\boldsymbol{\mathbf{D}}}_{\text{c}}\boldsymbol{\mathbf{\varphi}}_{\text{c}}+\boldsymbol{\mathbf{\varphi}}_{\text{s}}^{\intercal}\hat{\boldsymbol{\mathbf{D}}}_{\text{s}}\boldsymbol{\mathbf{\varphi}}_{\text{s}}$ (39)
where $\boldsymbol{\mathbf{\varphi}}_{\gamma}=[\varphi_{\gamma}(t^{+},x),\varphi_{\gamma}(t^{-},x)]^{\intercal}$, and the entries of the matrix $\hat{\boldsymbol{\mathbf{D}}}_{\gamma}$ ($\gamma=\text{c, s}$) are
$\hat{\boldsymbol{\mathbf{D}}}_{\gamma}^{\alpha,\beta}=\left[G_{\gamma}^{\alpha,\beta}\right]^{-1}$ (40)
@@
IV.3.2 Shot noise of the strongly correlated edge.
We now consider shot noise for the strongly correlated edge. The analysis is largely the same as for the weakly correlated edge, as the current is still carried by the $\varphi_{\text{c}}$ boson in the strongly correlated edge. The steps from Eqs. 42-48 are unchanged. However, instead of Eq. 49 we have (see Appendix F)
@@
In this work, we analyzed the role of interactions at the edge of the double quantum spin Hall insulators realized in twisted TMDs. We have shown that the edge has two symmetry preserving phases: a weakly correlated phase, a strongly correlated phase. The weakly correlated edge has two pairs of helical modes and is adiabatically connected to two decoupled quantum spin Hall edges. The strongly correlated edge has a single helical mode, and notably, is gapped to single electrons. In contrast, pairs of electrons remain gapless, leading the strongly correlated edge to have a Fano factor of $\mathcal{F}=2$ in shot noise measurements. Since $\mathcal{F}=1$ for the weakly correlated edge, shot noise measurements are a concrete method for distinguishing between the two possible edges.
While we have only considered the double quantum spin Hall insulator, it is straightforward to generalize these results to $N$-fold QSHIs. In this case, there is a non-trivial strongly correlated edge where there is a gap to single electrons, and the local degrees of freedom carry charge $Ne$. The Fano factor of the strongly correlated $N$-fold QSHI edge will therefore be $\mathcal{F}=N$. In particular, if the experimentally-observed triple quantum spin Hall insulator [19, 21] realizes a strongly correlated edge, the Fano factor will be $\mathcal{F}=3$.
In general, the existence of the strongly correlated edge depends on the Luttinger parameter of the edge theory. The Luttinger parameter, in turn, depends on the velocity of the edge modes with a lower bare velocity producing a smaller Luttinger parameter and stronger correlations. There are two main factors that determine the bare velocity of the edge modes: the properties of the bulk band dispersion and the microscopic details of the edge termination. The bulk bands of twisted TMDs are quite flat, making them potential candidates for realizing the strongly correlated DQSHI edge. However, the physical boundaries of the twisted TMD must still be taken into account to confirm this expectation. To measure the doubled Fano factor of the strongly correlated edge, it is also important that the constriction used to create the QPC does not destabilize the strongly correlated edge itself. If the single electron gap vanishes at the QPC, then individual electrons can tunnel between the edges, as in the weakly correlated edge. More sophisticated modeling of the boundaries and constrictions of a realistic TMD-device is needed to determine if if the effects discussed here will occur or not.
@@ Acknowledgments.
$\boldsymbol{\mathbf{\varphi}}^{\prime}\equiv\boldsymbol{\mathbf{W}}^{-1}\boldsymbol{\mathbf{\varphi}}$ (53)
@@
[…]
@@
$\boldsymbol{\mathbf{\varphi^{\prime}}}=\begin{pmatrix}\varphi_{+}\\ \varphi_{-}\\ \theta_{-}\\ \theta_{+}\end{pmatrix}$. (55)
@@
$\begin{split}\mathcal{L}&=-\frac{1}{4\pi}\partial_{x}(\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\varphi}}^{\prime})^{\intercal}\boldsymbol{\mathbf{K}}\partial_{t}(\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\varphi}}^{\prime})+\frac{1}{4\pi}\partial_{x}(\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\varphi}}^{\prime})^{\intercal}\boldsymbol{\mathbf{V}}\partial_{x}(\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\varphi}}^{\prime})\\ &=-\frac{1}{4\pi}\partial_{x}\boldsymbol{\mathbf{\varphi}}^{\prime\intercal}\underbrace{(\boldsymbol{\mathbf{W}}^{\intercal}\boldsymbol{\mathbf{K}}\boldsymbol{\mathbf{W}})}_{\mathclap{\boldsymbol{\mathbf{K}}^{\prime}}}\partial_{t}\boldsymbol{\mathbf{\varphi}}^{\prime}+\frac{1}{4\pi}\partial_{x}\boldsymbol{\mathbf{\varphi}}^{\prime\intercal}\underbrace{(\boldsymbol{\mathbf{W}}^{\intercal}\boldsymbol{\mathbf{V}}\boldsymbol{\mathbf{W}})}_{\mathclap{\boldsymbol{\mathbf{V}}^{\prime}}}\partial_{x}\boldsymbol{\mathbf{\varphi}}^{\prime}\end{split}$ (56)
where the transformed matrix $\boldsymbol{\mathbf{K}}^{\prime}$ is now
$\boldsymbol{\mathbf{K}}=\begin{pmatrix}0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\\ 1&0&0&0\\ \end{pmatrix}$ (57)
@@
$\boldsymbol{\mathbf{V}}^{\prime}=\boldsymbol{\mathbf{W}}^{\intercal}\boldsymbol{\mathbf{V}}\boldsymbol{\mathbf{W}}=\begin{pmatrix}v-v^{\prime}+u-u^{\prime}&&&\\ &v-v^{\prime}-u+u^{\prime}&&\\ &&v+v^{\prime}-u-u^{\prime}&\\ &&&v+v^{\prime}+u+u^{\prime}\end{pmatrix}$. (58)
@@
[…]
@@
$\boldsymbol{\mathbf{V}}=\begin{pmatrix}v_{1}&-v_{1}^{\prime}&u&-u^{\prime}\\ -v_{1}^{\prime}&v_{1}&-u^{\prime}&u\\ u&-u^{\prime}&v_{2}&-v_{2}^{\prime}\\ -u^{\prime}&u&-v_{2}^{\prime}&v_{2}\end{pmatrix}$ (61)
@@
where $w=v_{1}-v_{2}-v_{1}^{\prime}+v_{2}^{\prime}$ and $w^{\prime}=v_{1}-v_{2}+v_{1}^{\prime}-v_{2}^{\prime}$. However, we show in Appendix C that including these interactions do not affect the main results, since the propagators between the ++ and - sectors vanish in the limit as $\omega\to 0$.
@@
$\psi_{1\uparrow\text{L}}^{\dagger}\psi_{1\uparrow\text{R}}+\text{H.c.}$ $\sim\cos\left(\varphi_{\text{c}}+\varphi_{\text{n}+}+\varphi_{\text{s}+}+\varphi_{\text{n}-}\right)$ (64)
$\psi_{1\downarrow\text{L}}^{\dagger}\psi_{1\downarrow\text{R}}+\text{H.c.}$ $\sim\cos\left(\varphi_{\text{c}}+\varphi_{\text{n}+}-\varphi_{\text{s}+}-\varphi_{\text{n}-}\right)$ (65)
@@ Appendix C Low-energy asymptotics of the retarded propagator
In this section, we slightly generalize the calculation found in [67] to the case of two coupled inhomogeneous Luttinger liquids with unequal velocities and scattering parameters. We will obtain the low-energy asymptotic form of the various correlators, which we show are determined by the values of the parameters in the leads. We demonstrate that the original result of [67] still holds, and that the propagators between the ++ and - sectors vanish as $\omega\to 0$. This means that we may ignore this interaction term for the purposes of this analysis, which is concerned exclusively with the zero-frequency limit of all quantities.
@@
[…]
The entries of the matrix $\hat{\boldsymbol{\mathbf{D}}}$ are the following differential operators:
@@ C.3 Region II: $0<x<x^{\prime}$
By acting with the appropriate differential operators on the equations (for instance, $D^{\prime}$ on (88) and with $D_{-}$ on (90)), one finds that $_{\pm}, GG^{\prime}$ and $G^{\prime\prime}$ are all solutions to the same homogeneous, fourth-order, linear differential equation:
@@
$D_{-}D_{+}=\frac{v_{-}v_{+}}{K_{-}K_{+}}\partial_{x}^{4}-\omega^{2}\left(\frac{v_{+}^{2}+v_{-}^{2}}{v_{-}K_{-}v_{+}K_{+}}\right)+\frac{\omega^{4}}{v_{-}K_{-}v_{+}K_{+}}\\$ (93)